# Large Segmentation Design of Robust Photonic Quantum Gates

### v1

@@ Abstract
We analyze the fidelity and power loss of photonic quantum gates, implemented by a segmentation of coupled optical waveguides, as functions of the number of segments. We show that designs with a large number of segments mitigate effectively the systematic errors for a wide range of error correlation strength and variance. We quantify the design roughness effect on the power loss and analyze the performance of various methods for smoothing the configuration of the waveguides. We observe an asymptotic scaling behaviour of the gate fidelity and power loss with the number of segments.
@@ I Introduction
The quest for harnessing the potential of quantum phenomena has been a driving force behind intensive fundamental research and technology efforts, including Quantum Information Processing (QIP) bennett2000quantum .
Quantum computers hold the promise for an exponential speedup in solving certain complex computational problems and a capability to decrypt classical encryption protocols in real-time, thereby opening new frontiers in computing and cryptography. Despite substantial theoretical advancements in quantum algorithms QuantumAlgorithms1 ; QuantumAlgorithmSurvey , the practical realization of a quantum information processor remains a formidable challenge.
One of the paramount obstacles in achieving practical quantum computation lies in the noise and errors that affect the performance of the fundamental building blocks of quantum algorithms: the qubit’s state preparation, the quantum gates, and the measurement process Nielsen2000 . These errors and noise have various sources, such as fabrication imperfections Kyoseva2019 , inaccurate coupling strengths and the interaction with the classical environment noiseoneoverf . In response to these challenges, two primary routes have been taken: error-correcting codes QEC1 and error mitigation schemes NoiseMitigation2 ; QNoiseMitigation1 .
Error-correcting codes construct logical qubits using additional entangled physical qubits and can significantly outperform their physical counterparts. Notable among these codes are stabilizer codes, exemplified by the surface code Fowler2012 , which exhibits a relative tolerance to local errors. Fault-tolerant quantum computation using error correcting codes requires the fidelity of the physical quantum gates to be above a certain threshold, making the reduction of the physical errors to be of utmost importance NoiseThreholdQC ; QuantumErrorCorrection .
Error mitigation schemes seek to reduce errors using sophisticated implementation and techniques, for a review see QNoiseMitigation1 ; NoiseMitigation2 . One such framework is the segmented composite pulse design, which was originally used in atomic physics and nuclear magnetic resonance, employing a combination of constant pulses to minimize errors during the system evolution CPNMR ; CPNMR2 ; CPAtomicPhysics ; CPAtomicPhysics2 . This error mitigation technique holds a promise for improving the performance of complex quantum algorithms, and has a broad applicability in diverse quantum computing realizations of quantum information processing devices.
In SmoothCP smooth designs were proposed in a general composite pulses scheme, and HUSAIN2013145 suggested symmetric and antisymmetric composite pulses with a large number of pulses in NMR. It has been recently shown in Kaplan2023 that this method can be implemented in Silicon photonics KLM for designing robust high fidelity quantum gates by varying the waveguide geometry, and specifically its cross-section. The focus in Kaplan2023 has been on having a small number of segments, under the assumption of fully correlated noise of the different waveguides. It was proven in YaronIdoNoiseThreshold that there is a correlation strength threshold for the segmented design to be an effective error mitigating scheme.
The aim of the current work is to analyze the composite pulse scheme for optimizing robust designs with a large number of segments. We consider both the fidelity of the quantum gates as well as the power loss of the designs for varying systematic noise correlations. Applied to the photonic dual-rail realization, our methods yield robust high-fidelity unitary designs for single-qubit and hence also for the probablistic double-qubit gates such as CZ and CNOT Knill2001 . We consider different manufacturing configurations in Silicon and Lithium Niobate, as Silicon is widely used in this field, while the Lithium Niobate is considered to be a promising material for designing active QIPs.
Our main contributions are :
  1. 1.
    Enhanced Quantum Gate Robustness: We demonstrate numerically that increasing the number of composite segments $N$ results in a substantial improvement in the quantum gate mean fidelity $\mathbf{E}\left[F\right]$, and a reduction in its standard deviation $\sigma_{F}\left[F\right]$ with a certain scaling with $N$. We further show that $N\sim O(10)$ is effective for most practical cases.
  2. 2.
    Power Loss Analysis: We quantify the impact of increasing $N$ on the power loss in the quantum gate due to the back-scattering effects in the segments. We argue analytically and show numerically that the power loss grows linearly with the number of segments, which makes it necessary to apply smoothing methods to the large $N$ designs.
  3. 3.
    Correlation Strength: We study the impact of error correlation strength between the different segments on the gate fidelity. We show that enhancing the correlation strength among the errors in different segments leads to a significant increase in the expected fidelity of the quantum gate, hence providing valuable insights to strategies for choosing the optimal number of segments given a correlation strength. For a given required gate robustness, larger correlation strength implies that less segments are needed.
  4. 4.
    Adiabatic and Fresnel Designs: We introduce smooth segmented gates designs based on the principles of the Adiabatic and Fresnel scattering, that exhibit a reduced power loss while keeping the same fidelity performance. This offers a complementary tool to the traditional segmentation method.
The paper is organized as follows: In Section II, we briefly review the composite pulse design of quantum gates and the error model in the optical waveguides framework. In Section III we present the ingredients of the cost function optimization framework including the optimization range, the quantum gate fidelity and the smooth design conditions for minimizing the power loss. In Section IV we show the optimization results for the gates fidelity and power loss for a varying number of segments and for different smooth designs. Section V is devoted to a discussion and conclusions. Additional details are provided in four appendices.
II Segmented Composite Pulses
@@
Figure 1: We divide each of the two waveguides to $N$ segments. Each segment $i$ is characterized by its coupling $\kappa_{i}$, mode-mismatch coefficient $\Delta_{i}$ and length $z_{i}$. The interpolation map for $\kappa$ and $\Delta$ as a function of the waveguides widths $W_{1},W_{2}$, in $nm$ units. These results are extracted from a Lumerical FDTD Eigensolver simulation, for Silicon on insulator waveguides platform with a height of $220nm$ and a gap of $1\mu m$. One of the waveguides was illuminated with $1.55\mu m$ light source.
II.1 Robust Composite Quantum Gates
@@
The presence of noise within the physical two-level system leads to deviations from the desired quantum gate operation. To quantify the extent to which the noisy gate deviates from the ideal gate, we define the gate fidelity $F$ by the trace norm:
@@
where $U_{\text{ideal}}$ represents the desired ideal unitary gate, and $U(\boldsymbol{\epsilon})$ represents its physical realization, which depends on a set of $m$ jointly distributed random errors denoted as $\boldsymbol{\epsilon}=\{\epsilon^{a}\}_{a=1}^{m}$. The fidelity is a random function that takes values in the interval range $[0,1]$, where zero and one correspond to the absence of errors and to the maximal deviation from the desired quantum gate operation, respectively. The primary objective of our analysis is to maximize the mean fidelity, denoted as $\mathbf{E}\left[F\right]$, and to minimize its standard deviation, denoted as $\sigma[F]$:
@@
The composite pulses is a framework for error mitigation where $U(\boldsymbol{\epsilon})$ is realized as a product of $N$ physical unitary gates:
where $\vec{\boldsymbol{\epsilon}}=(\boldsymbol{\epsilon}_{1},...,\boldsymbol{\epsilon}_{N})$, $\boldsymbol{\epsilon}_{i}=\{\epsilon^{a}_{i}\}_{a=1}^{m}$, and $U_{i}$ are constructed to achieve higher fidelity. We will consider this error mitigation scheme in the framework of dual-rail waveguides in integrated photonics.
II.2 Coupled Optical Waveguides
Consider the following physical photonic realization of single-qubit quantum gates, which is also the basis for constructing probabilistic two-qubit gates AmnonYariv . A coupled system of two waveguides can be described by two parameters: the coupling coefficient $\kappa$ and the mode-mismatch coefficient $\Delta$ (known also as the detuning coefficient). The coupling coefficient is given by the integral over the spatial modes-overlap:
[…]
where $(1)$ and $(2)$ refer to the two waveguides in the dual-rail representation, $(x,y)$ are the transverse directions, $z$ denotes the propagation direction and $\epsilon(x,y)$ and $\epsilon^{(2)}(x,y)$ are the permittivity distributions of the coupled waveguides system and of waveguide $(2)$ alone, respectively. The mode-mismatch coefficient is determined by the difference between the propagation constants of the two waveguides:
$\Delta=\frac{\beta_{1}-\beta_{2}}{2}\equiv\frac{2\pi}{\lambda}\left(n_{{eff}_{1}}-n_{{eff}_{2}}\right)\ ,$ (5)
where $n_{{eff}_{i}}$ are the effective refraction indices in the first and second waveguides, respectively, and depend on the geometry of the waveguides. A simple solution for the propagation differential equation is constant coupling and mode-mismatch, which we refer to as one segment. It is given by (see appendix A):
$+i\frac{\Delta}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{Z}\ ,$ (6)
where $\Omega_{g}\equiv\sqrt{\Delta^{2}+\kappa^{2}}$, and $\mathbf{I},\mathbf{X},\mathbf{Z}$ are the identity matrix and Pauli matrices, respectively.
II.3 Segmentation
We will analyze quantum gates designs that consist of $N$ segments (see Fig. 1). Supported by experimental evidence ExperimentCP , we consider the case where in each segment, the noise affects both waveguides identically (full correlation), i.e. $\epsilon^{(1)}_{k}=\epsilon^{(2)}_{k},~{}k=1,...,N$. This implies a possible significant error in the coupling coefficients $\kappa_{k}$, but affects less the mode-mismatch coefficients $\Delta_{k}$. As to the correlation between the errors of the different segments, we will initially assume a full correlation between them, and will later relax this condition and analyze the case of a general correlation matrix. The $N$-segmented gate reads:
$U^{(N)}=\prod_{k=1}^{N}U(\kappa_{k},\Delta_{k},z_{k},\epsilon_{k})\ ,$ (7)
where $z_{k},\kappa_{k},\Delta_{k},\epsilon_{k}$ are the length of the $k^{th}$-segment, its coupling, its mode-mismatch coefficient and its noise, respectively.
II.4 Power Loss
Implementation of the composite pulse error mitigation scheme in Silicon photonics or in any other physical platform is accompanied by power loss due to back-scattering. This back-scattering is a direct consequence of the non-smooth profile of the segmented waveguide’s width, and captures the total power loss (see Appendix B).The gate reflection power loss is total power transmission coefficient $P_{t}$:
@@ II.4 Power Loss
Specifically, in photonics, we can estimate the back-scattering by the Fresnel equation that governs the reflection and transmission coefficients of a light transmitted between two different materials with refractive indices $n_{1}$ and $n_{2}$ (we consider the intensity ):
$R=\left(\frac{n_{1}cos\theta_{i}-n_{2}cos\theta_{t}}{n_{1}cos\theta_{i}+n_{2}cos\theta_{t}}\right)^{2}\ ,$ (9)
where we chose TE polarization and $\theta_{i}$ and $\theta_{t}$ are the incident and transmitted light angle, respectively. In guided propagation, the refractive index is replaced by $n_{eff}$, which is calculated from the overlap integral of the spatial cross-section field profile with the cross-section refractive index.
For our analysis, focusing on the TE polarization (the basic mode) and assuming small angles related to the $z$ axis, we evaluate the total total reflection coefficient as:
$R=1-T=1-\prod_{i=1}^{N-1}\left(1-\left(\frac{n_{eff_{i+1}}-n_{eff_{i}}}{n_{eff_{i+1}}+n_{eff_{i}}}\right)^{2}\right)\ .$ (10)
III Optimization Framework
Our optimization search method is subject to several constraints. While our goal is to obtain high fidelity quantum gates, we aim to achieve it with as smooth as possible design in order to minimize the power loss. As discussed, abrupt changes in the waveguide’s width increase the power loss due to back scattering, and also give rise to higher spatial modes in the waveguide that distort the signal, hence changing the qubit state. Moreover, waveguide designs must be physically implementable subject to the manufacturing resolution.
We divide each waveguide into a sequence of $N$ small segments, each of length $z_{i}$ as shown in Fig. 1. The two waveguides are separated by a gap (defined as the distance between the centers of the waveguides) and have a height $h$. Our optimization uses the Lumerical FDTD results as a look-up table to inverse-map the geometry to the coupling coefficient and the mode-mismatch parameters.
III.1 Fidelity
We will assume that the errors in the two waveguides widths $W_{1}$ and $W_{2}$ are fully correlated and can be sampled from a Normal distribution with zero mean and a standard deviation $\sigma$, $D\sim\mathcal{N}(0,\sigma)$. For our analysis, we consider the dimensionless parameter $\sigma_{\%}\equiv 100\frac{\sigma}{W_{mean}}$, where

### v2

# Robust photonic quantum gates with large number of waveguide segments

@@ Abstract
Realizing quantum information processors is challenged by errors and noise across all platforms. While composite segmentation schemes have been proposed in many systems, their application to photonic quantum gates in dual-rail configurations has only recently been demonstrated. However, prior research has been limited to a small number of segments, full noise correlation, and has overlooked the inherent power loss in such designs. Here, we study the fidelity and power loss of composite designs for photonic quantum gates with a high number of segments of varying geometrical widths. Using numerical simulations, we analyze the relationship between gate performance and the number of waveguide segments, accounting for statistical error correlations and variances. Beyond effectively reducing the errors, an asymptotic scaling pattern of quantum gate fidelity and power loss is observed as the number of segments increases. This analysis is examined in Silicon and Lithium Niobate platforms, addressing practical implementation challenges. Our findings demonstrate that optimized multi-segment waveguide geometrical designs significantly enhance the robustness and efficiency of photonic quantum gates, paving the way for more reliable quantum information processors.
@@ I Introduction
The realization of a practical quantum information processor remains a great challenge, as errors and noise inevitably arise during the implementation of its core elements. These challenges are universal across various physical platforms for quantum information processing bennett2000quantum ; Nielsen2000 , including superconducting circuits Kjaergaard2020 , trapped ions Bruzewicz2019 , and photonic systems Flamini2018 ; giordani2023integrated . Among these, integrated photonics, particularly dual-rail waveguide platforms, has emerged as a promising approach due to its scalability and compatibility with existing fabrication technologies.
In dual-rail photonic systems, errors primarily stem from fabrication imperfections Kyoseva2019 , inaccurate coupling strengths, and environmental interactions noiseoneoverf , which compromise the fidelity of the quantum state preparation and quantum gates, the fundamental building blocks of any quantum algorithm Nielsen2000 . For example, deviations in waveguide geometries from the intended design, such as variations in waveguide width, can significantly impact the performance of dual-rail single-photon and two-photon gates. These deviations induce Pauli errors, which in turn lead to a reduction in gate fidelity melati2014real . Various strategies have been developed to address these challenges QNoiseMitigation1 ; NoiseMitigation2 . Error correction codes, for example, construct logical qubits from entangled physical qubits to reduce errors QEC1 . However, these methods impose stringent fidelity thresholds on the underlying physical gates, requiring a fidelity error of less than $1\%$ in each of the physical photonic gates NoiseThreholdQC ; QuantumErrorCorrection ; shi2022high ; noiri2022fast .
One promising error mitigation approach, which reduces physical errors through advanced design and implementation techniques, is the composite design method, originally developed in atomic physics and nuclear magnetic resonance (NMR) CPNMR ; CPNMR2 ; CPAtomicPhysics ; CPAtomicPhysics2 . This method employs systematic parameter variations to counteract errors during system evolution. In NMR, large number of pulses were suggested, using symmetric and antisymmetric composite pulses designs HUSAIN2013145 , and a smooth composite pulses scheme have been proposed in Ref. SmoothCP .
Recent work has adapted this approach to integrated photonics, demonstrating the potential for high-fidelity dual-rail quantum gates by segmenting waveguides and varying their geometries Kyoseva2019 ; Kaplan2023 . However, prior studies primarily considered designs with a small number of segments and assumed fully correlated fabrication errors Kaplan2023 . Subsequent findings indicate that effective error mitigation requires the correlation strength to remain below a specific threshold YaronIdoNoiseThreshold . Moreover, earlier analyses did not evaluate the power loss associated with segmentation, an essential factor for practical implementation.
Figure 1: Comparison of composite segment design with variable number of segments. (a) Two coupled waveguides, with widths denoted as $W_{0}$ and $W_{1}$, a gap between them and a defined length. (b) Optical waveguide gates in dual-rail encoding for $X$ (Pauli), gate, along with its graphical representation on the Bloch sphere. (c) Fabrication noise model with spatial correlation between segments, for a small and large number of segments. (d) Comparison of Fidelity and Power Loss between a small and large number of segments. Increasing either the number of segments or the noise spatial correlation improves Fidelity. Our proposed smooth design significantly reduces power loss, in addition to increasing Fidelity.
Here, we study the fidelity and power loss of photonic quantum gates composed of increased number of waveguides’ segmentations. We find that as the number of segments increases, the errors associated with variations in waveguide geometrical widths are significantly mitigated, leading to substantially higher gate fidelity. Using numerical simulations on Silicon-on-Insulator (SOI) and Thin-Film Lithium Niobate (TFLN) platforms, we demonstrate that increasing the number of waveguide segments $N$, not only improves the average quantum gate fidelity $\mathbf{E}\left[F\right]$ but also reduces its standard deviation $\sigma_{F}\left[F\right]$, with a specific scaling with $N$.
Furthermore, we analyze the impact of segmentation on power loss due to back-scattering effects. Our analytical and numerical results show that the power loss increases linearly with the number of segments, underscoring the necessity of employing smoothing techniques for designs with large $N$. To address this, we propose smooth segmented photonic quantum gate designs based on adiabatic and Fresnel reflection principles. These designs effectively reduce power loss while maintaining high fidelity performance. Finally, we conclude that a segmentation number $N\sim O(10)$ is sufficient for most practical implementations, providing a balance between fidelity enhancement and manageable power loss. This work advances the development of practical quantum dual-rail gates for photonic integrated platforms, and quantum information processors in integrated photonics.
The paper is organized as follows: Section II introduces the theoretical background, including dual-rail encoding within the framework of integrated optics, as well as the necessary background on composite schemes. In Section III, we formulate the optimization problem and define the cost function, incorporating the optimization range, quantum gate fidelity, and smooth design conditions aimed at minimizing power loss. Section IV presents the optimization results for gate fidelity and power loss across varying numbers of segments and different smooth design configurations. Finally, Section V provides a summary and outlook. Additional technical details are included in four appendices.
II Theoretical background
II.1 Dual-rail realization for of two-mode dynamics
In the integrated dual-rail realization, two closely spaced optical waveguides form a coupled system that emulates the dynamics of a quantum two-level system AmnonYariv . This system, governed by coupled-mode theory (CMT), is described by a unitary propagator that captures its evolution over the propagation length:
$i\partial_{z}\left(\begin{array}[]{l}E_{0}(z)\\ E_{1}(z)\end{array}\right)=\left(\begin{array}[]{cc}-\Delta(z)&\kappa^{*}(z)\\ \kappa(z)&\Delta(z)\end{array}\right)\left(\begin{array}[]{l}E_{0}(z)\\ E_{1}(z)\end{array}\right)\ ,$ (7)
where $E_{0}(z)$ and $E_{1}(z)$ are the electrical fields at the of being in states $|0\rangle$ and $|1\rangle$, the mode mismatch $\Delta(z)$ arises from differences in the effective refractive indices, $n_{\text{eff},1}$ and $n_{\text{eff},2}$ of the waveguides:
$\Delta(z)=\frac{\beta_{1}(z)-\beta_{2}(z)}{2}\equiv\frac{2\pi}{\lambda}\left(n_{{eff}_{1}}(z)-n_{{eff}_{2}}(z)\right)\ ,$ (8)
where $\beta_{1}$ and $\beta_{2}$ are the propagation constants of the two waveguides.
The coupling coefficient $\kappa$, resulting from the evanescent field overlap between the waveguide modes, is defined as:
$\kappa(z)=\frac{c}{4\pi\lambda}\iint$ $\left[\epsilon(x,y,z)-\epsilon^{(1)}(x,y,z)\right]\cdot$
[…]
where $(0)$ and $(1)$ refer to the two waveguides, $(x,y)$ are the transverse directions, $z$ denotes the propagation direction, $\vec{E}_{\perp}^{(0)}(x,y,z)$ and $\vec{E}_{\perp}^{(1)}(x,y,z)$ are the transverse electric field modes for each of the waveguides, and $\epsilon(x,y,z)$ and $\epsilon^{(1)}(x,y,z)$ are the permittivity distributions of the coupled waveguides system (namely, the profile of the two waveguides together) and of waveguide $(1)$ alone, respectively AmnonYariv . $c$ is the speed of light and $\lambda$ is the wavelength. A diagram of two coupled waveguides is shown in Fig. 1 a.
The unitary propagator of such a system is:
$U(z,0)=\mathcal{Z}\left[\exp\left[-i\int_{0}^{z}\left(\begin{array}[]{cc}-\Delta\left(z^{\prime}\right)&\kappa^{*}\left(z^{\prime}\right)\\ \kappa\left(z^{\prime}\right)&\Delta\left(z^{\prime}\right)\end{array}\right)\right]dz^{\prime}\right]\ ,$ (12)
where $\mathcal{Z}$ denotes z-ordering. In cases where $\kappa$ and $\Delta$ are independent of z, the propagator simplifies to:
$U(z,0)=\exp\left[-iz\left(\begin{array}[]{cc}-\Delta&\kappa^{*}\\ \kappa&\Delta\end{array}\right)\right]$ (15)
Alternatively, using the well-known relation $e^{-i\alpha\vec{\sigma}\cdot\hat{n}}=cos(\alpha)\mathbf{I}+\vec{\sigma}\cdot\hat{n}sin(\alpha)$, for a unit vector $\hat{n}$, and $\vec{\sigma}=(\mathbf{X},\mathbf{Y},\mathbf{Z})$ a vector of Pauli matrices, the unitary operator can be written as:
$+i\frac{\Delta}{\Omega_{g}}\sin\left(\Omega_{g}z\right)\mathbf{Z}\ ,$ (16)
while we use use the transfer frequency determined by the effective coupling parameter $\Omega_{g}=\sqrt{\kappa^{2}+\Delta^{2}}$.
The dual-rail encoding in two coupled optical waveguides forms a quantum two-level system, where light propagation emulates $SU(2)$ dynamics, and is elegantly represented on the Bloch sphere (Fig. 1 (b)), where the evolution of light in the coupled waveguides mirrors Rabi oscillations in atomic physics. For symmetric configurations ($\Delta=0$), the system achieves Complete Power Transfer (CPT), between the two levels. This $SU(2)$ framework underpins the implementation of integrated photonic quantum gates and sets the stage for advanced multi-qubit operations via concatenated or segmented couplers.
Realizing two-qubit gates, such as the Controlled-NOT (CNOT) or Controlled-Z (CZ) gates, typically requires additional ancilla modes, measurements, and linear optical elements, and these gates often exhibit a non-deterministic nature KLM . A well-known example is the CNOT gate, which can be implemented using beam splitters, phase shifters, and single-photon detectors Ralph2002 .
By carefully arranging these components, conditional operations are performed, where the state of one qubit (the control qubit) governs the state of another (the target qubit). The universality of quantum computing in the dual-rail encoding scheme is achieved through the combination of single-qubit and two-qubit gates. This universal gate set enables the construction of arbitrary quantum circuits, paving the way for the implementation of complex quantum algorithms.
II.2 Coherent Noise Analysis
@@
Figure 2: $N$ segments composite configuration. We divide each of the two waveguides to $N$ segments. Each segment $i$ is characterized by its coupling $\kappa_{i}$, mode-mismatch coefficient $\Delta_{i}$ and length $z_{i}$. The interpolation map for $\kappa$ and $\Delta$ as a function of the waveguides widths $W_{0},W_{1}$, in $nm$ units. These results are extracted from a Lumerical FDTD Eigensolver simulation, with geometry resolution of $5nm$, and a bilinear interpolation to evaluate the values in-between.
Coherent noise in photonic quantum gates, caused by fabrication imperfections, introduces systematic errors in the form of random unitary rotations. In the photonic regime, these errors are primarily attributed to variations in the waveguide widths during the fabrication process, which significantly impact the coupling coefficient $\kappa$ and mode mismatch $\Delta$. Additionally, due to the nature of errors arising during the fabrication of coupled waveguides, these errors are fully correlated between the two coupled waveguides, further compounding their effect on gate fidelity ExperimentCP . Such systematic deviations degrade gate performance and accumulate in multi-qubit operations, necessitating mitigation to meet the stringent thresholds required for quantum error correction and measurement-based quantum computation Nielsen2000 .
@@
To quantify the extent to which the noisy gate deviates from the ideal gate, we define the gate fidelity $F$ by the trace norm:
@@
where $U_{\text{ideal}}$ represents the desired ideal unitary gate, and $U(\boldsymbol{\epsilon})$ represents its physical realization, which depends on a set of $m$ jointly distributed random errors denoted as $\boldsymbol{\epsilon}=\{\epsilon_{a}\}_{a=1}^{m}$. The fidelity is a random function that takes values in the interval range $[0,1]$, where zero and one correspond to the absence of errors and to the maximal deviation from the desired quantum gate operation, respectively.
The primary objective of our analysis in the context of error mitigation, is to maximize the mean fidelity, denoted as $\mathbf{E}\left[F\right]$, and to minimize its standard deviation, denoted as $\sigma[F]$:
@@
Implementation of the composite pulse error mitigation scheme in integrated photonics, or in any other physical platform, is accompanied by power loss due to several reasons such as absorption or scattering. The back-scattering is a direct consequence of the non-smooth profile of the segmented waveguides width profiles, and captures the total power loss (see Appendix B). The gate reflection power loss is total power transmission coefficient $P_{t}$:
@@
Due to the changes in the effective dispersion coefficients between the segments of the waveguides, we estimate the back-scattering by the Fresnel equation, with effective refractive indices of the two consecutive segments:
$T_{k}=1-\left(\frac{n_{\text{eff},k}-n_{\text{eff},k+1}}{n_{\text{eff},k}+n_{\text{eff},k+1}}\right)^{2}\ ,$ (20)
The total reflection coefficient is given by:
$R=1-P_{t}\ .$ (21)
II.3 The Composite Segmentation Method
The composite pulses method is well-known for its ability to mitigate errors by decomposing a single gate into a sequence of smaller unitary operations. By strategically leveraging the correlation between error factors, this sequence can be meticulously designed to cancel out errors effectively. In the dual-rail implementation, as demonstrated in Ref. Kaplan2023 , a two-waveguide coupler is divided into multiple sections with varying widths and shorter propagation lengths.
We define the gate $U(\boldsymbol{\epsilon})$ as a sequence of $N$ segments:
Where $\boldsymbol{\epsilon}_{i}$ are the error parameters (in our context, in the width of the waveguides), a single one for each gate. The $U_{i}$ are characterized by a its own $\kappa_{i},\Delta_{i},z_{i}$ parameters (coupling coefficient, mode mismatch, and length, respectively).
Therefore, designing a robust unitary gate $U_{cp}$ is actually finding the set of geometrical parameters (width, length, refractive index etc.), that determines $\{\kappa_{i},\Delta_{i},z_{i}\}_{i=1}^{3}$ (total of $3N$ degrees of freedom) such that the gate will be more robust to the presence of noise, namely, the fidelity between $U_{cp}$ and $U$ will be as close as possible to 1.
III High fidelity quantum gates using Large N segmentation
In this study, we focus on dual-rail waveguide platforms, presenting a detailed study of segmented and smooth waveguide designs of robust, high-performance quantum gates designs. Fig. 1 provides an overview of the work, including the key results from this paper: showcasing the relationship between fidelity and power loss against the number of segments (see Fig. 1c), as well as the impact of smoothing and noise correlation (see Fig. 1d).
Our aim is to obtain high-fidelity quantum gates with as smooth as possible design that will minimize the power loss, for any arbitrary number of segments $N$. As discussed, abrupt changes in the waveguides widths increase the power loss due to back scattering, and also give rise to higher spatial modes in the waveguide that distort the signal, hence changing the qubit state. Moreover, waveguide designs must be physically implementable subject to the manufacturing resolution.
We divide each waveguide into a sequence of $N$ small segments, each of length $z_{i}$ as shown in Fig. 2. The two waveguides are separated by a gap, defined as the distance between the centers of the waveguides, and have a height $h$. The optimization process involves three stages. Firstly, we construct a look-up table with the Lumerical FDTD simulator to map geometries to the coupling coefficient and mode-mismatch parameters inversely. Next, we establish the optimization cost function, including regularizes. Finally, we filter-out devices that are not physically realizable. The subsequent subsections will cover these stages in depth.
III.1 The Fidelity Term
As discussed, in our analysis we assume that the errors in the two waveguides widths $W_{0}$ and $W_{1}$ are fully correlated, and can be sampled from a Normal distribution with zero mean and a standard deviation $\sigma$, $D\sim\mathcal{N}(0,\sigma)$. For our analysis, we consider the dimensionless parameter $\sigma_{\%}\equiv 100\frac{\sigma}{W_{mean}}$, where